Mills Manufacturing Inc. had sales for January 2020 of $900,000. The units sold were 18,000, and the expenses for January 2020 were as follows:

Variable Fixed
Cost of goods sold $181,800 $115,200
Selling expenses 53,100 27,900
Administrative expenses 25,200 29,700

Please find contribution margin per unit and contribution margin ratio.
Answer #1

Total Variable Cost = Cost of Goods sold + Selling Expenses + Administrative Expenses

Total Variable Cost = $181,800+53,100+25,200 = $260,100

Total Sales Revenue = $ 900,000

Contribution Margin = Sales Revenue - Total Variable Cost = 900,000-260,100 = $639,900

1) Contribution Margin Per Unit = Contribution Margin / Units Sold

= $639,900 / 18000

= $35.55 (Answer)

2) Contribution Margin Ratio = Contribution Margin / Revenue

= 639,900 /900,000

= 71.10 % (Answer)